OpenStudy (tmagloire1):

Water is drained out of tank, shaped as an inverted right circular cone that has a radius of 16cm and a height of 8cm, at the rate of 4 cm3/min. At what rate is the depth of the water changing at the instant when the water in the tank is 6 cm deep? Show all work and include units in your answer. Give an exact answer showing all work and include units in your answer.

OpenStudy (tmagloire1):

ap calc ab problem

Vocaloid (vocaloid):

start with the volume equation V = (1/3)pi*r^2*h

Vocaloid (vocaloid):

then we want to try to eliminate r by using the proportion if r = 16 and h = 8, we know that r = 2h so we get V = (1/3)*pi*(2h)^2*h = (1/3)*pi*(4)h^3

Vocaloid (vocaloid):

then we differentiate w/ respect to time

Vocaloid (vocaloid):

V = (4)(1/3)*(pi)h^3 dV/dt = (4/3)(pi)(3)(h^2) * dh/dt

Vocaloid (vocaloid):

then it's just plug and chug from there

OpenStudy (tmagloire1):

So the ultimate goal after you plug in is to find dh/dt?

Vocaloid (vocaloid):

yup! dh/dt is the rate at which the depth/height is changing

Vocaloid (vocaloid):

we already know dV/dt (4) and h

OpenStudy (tmagloire1):

Wait when did we find h?

Vocaloid (vocaloid):

"At what rate is the depth of the water changing at the instant when the water in the tank is 6 cm deep?"

Vocaloid (vocaloid):

h = 6

OpenStudy (tmagloire1):

Ohhh ok my bad. so then this is what I got: 4 = (4/3)(pi)(3)(6^2) * dh/dt 4= 452.16 * dh/dt dh/dt = .0096 this sounds very wrong

Vocaloid (vocaloid):

well, it makes sense since the cone is very wide compared to its height

Vocaloid (vocaloid):

let me just check the numbers real quick

OpenStudy (tmagloire1):

actually i got .0088

Vocaloid (vocaloid):

@DanJS @Directrix would you mind checking my work? might have made a mistake but I don't see where

zepdrix (zepdrix):

`Give an exact answer showing all work and include units in your answer.` 0.0088 sounds correct! :) But that is an approximation. Oh, and maybe throw a negative in front to show that the water is height is decreasing.\[\large\rm h'=-\frac{1}{36\pi}\frac{cm}{\min}\]

OpenStudy (tmagloire1):

Oh okay it was so small it scared me for a second haha! Thanks for your help both of you!